# Quantum Bayes’ rule and Petz transpose map from the minimal change principle

### v2

@@ Abstract
satisfy the condition that the partial derivative of $\mathcal{L}$ is zero in every direction. Therefore, $C_{\mathcal{R}}$ is a local optimum of the program. Finally, due to the strict concavity of fidelity under the theorem’s hypotheses (proven as a separate lemma in the Supplemental Material), we conclude that $C_{\mathcal{R}}$ is the unique global maximum.
Discussion.—In this work, we have generalized the minimal change principle to the quantum case, offering a new formulation of the quantum Bayes’ rule. In particular, when fidelity is used as a figure of merit, the Petz transpose map, which is often only pretty good but suboptimal, has naturally emerged as the unique optimal solution in many relevant cases, confirming the central role of the Petz transpose map as a quantum Bayes’ rule. The consistency between the theory of statistical sufficiency (in which the Petz transpose map plays a central role) and the variational principle of minimal change suggests a wide range of applicability for the minimal change principle in all areas where the Petz transpose map has appeared to play a role, such as quantum information theory, quantum statistical mechanics, and many-body physics.
Further, by imposing additional constraints to the program Eq. 11, one could restrict the reverse process to a desired subset. We have solved the program Eq. 11 analytically, while for general cases, convex additional constraints maintains the convexity of the optimization program, for which efficient numerical algorithms can be adopted [43, 44]. With this approach, the minimal change principle may be extended to quantum combs [45, 46], quantum supermaps [47, 48], and quantum Bayesian networks [49, 15, 50, 51, 52], offering new belief update rules for them. The tools introduced in this work may also pave the way towards a fully quantum generalisation of entropy production and fluctuation theorems [53, 54].
Acknowledgments.—G. B. and V. S. acknowledge support from the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; and from the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). F. B. acknowledges support from MEXT Quantum Leap Flagship Program (MEXT QLEAP) Grant No. JPMXS0120319794, from MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe,” No. 21H05183, and from JSPS KAKENHI Grants No. 20K03746 and No. 23K03230.
@@ Appendix A Proof of Theorem 1
We use the definition of $Q_{\rm fwd}$ and $Q_{\rm rev}$ in Eqs. 6 and 8. Since we are first proving the existence of a local optimal solution of Eq. 11 in the form Eq. 12, it suffices to restrict $C_{\mathcal{R}}$ to be full-rank, and then $Q_{\rm rev}$ is full-rank. Recall that the variable we optimize over is $\mathcal{R}$, or equivalently the transpose of its Choi operator $C_{\mathcal{R}}^{T}$.
@@
$(\sqrt{\tau}\otimes\mathds{1}_{A})\Delta(\sqrt{\tau}\otimes\mathds{1}_{A})+\Lambda^{T}\otimes\mathds{1}_{A}$ $=0$ (38)
$\mathrm{Tr}_{A}[C_{\mathcal{R}}]-\mathds{1}_{B}$ $=0$ (39)
@@
where $\Delta_{B}:=-\tau^{-1/2}\Lambda^{T}\tau^{-1/2}$. $\Delta_{B}>0$ since $\Delta>0$. By Eq. 39, one gets $\mathds{1}_{B}=\mathrm{Tr}_{A}[C_{\mathcal{R}}^{T}]$. Substituting Eqs. 35 and 40 into this, one gets
@@
where $D:=\sqrt{\tau}\left(\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau}\right)^{-1/2}$. The above Choi operator corresponds to the map in Eq. 12. To see this, assuming the Kraus decomposition of $\mathcal{E}$ is $\mathcal{E}(\rho)=\sum_{k}K_{k}\rho K_{k}^{\dagger}$, one computes $C_{\mathcal{R}}^{T}$ from Eq. 12 by definition of the Choi operator as
@@
By Eq. 22, $\Delta=\sqrt{\rho}\left(\sqrt{\rho}\,\sigma\sqrt{\rho}\right)^{-1/2}\sqrt{\rho}=\sqrt{\rho}M^{-1}\sqrt{\rho}$. Using Eqs. 32 and 63, its differential is
@@
Combining this with Eq. 62, we get the second-order derivate of $f$

### v3

@@ Abstract
satisfy the condition that the partial derivative of $\mathcal{L}$ is zero in every direction. Therefore, $C_{\mathcal{R}}$ is a local optimum of the program. Finally, due to the strict concavity of fidelity under the theorem’s hypotheses (proven as a separate lemma in Appendix A), we conclude that $C_{\mathcal{R}}$ is the unique global maximum.
Discussion.—In this work, we have generalized the minimum change principle to the quantum case, offering a new formulation of the quantum Bayes’ rule. In particular, when fidelity is used as a figure of merit, the Petz transpose map, which is often only pretty good but suboptimal [52, 53], has naturally emerged as the unique optimal solution in many relevant cases, confirming the central role of the Petz transpose map as a quantum Bayes’ rule. The consistency between the theory of statistical sufficiency (in which the Petz transpose map plays a central role) and the variational principle of minimum change suggests a wide range of applicability for that principle in all areas where the Petz transpose map has appeared to play a role, such as quantum information theory, quantum statistical mechanics, and many-body physics.
Besides fidelity, one may consider other divergences as the figure of merit, such as the Hilbert–Schmidt distance [54] or the Umegaki [55, 56] or Belavkin–Staszewski [57, 58] quantum relative entropies, and wonder whether they lead to the same quantum Bayes’ rule. In the special case of $\tau=\mathcal{E}(\gamma)$, they always do, because the Petz transpose map produces $Q_{\rm fwd}=Q_{\rm ref}$ and any reasonable divergence should reach optimum for two identical operators. However, in the general case $\tau\neq\mathcal{E}(\gamma)$, even if $[\tau,\mathcal{E}(\gamma)]=0$, numerical optimization shows that the optimum points may differ for different divergences. It is thus interesting to explore whether different quantum Bayes’ rules will arise from those various choices.
Further, by imposing additional constraints to the program Eq. 15, one could restrict the reverse process to a desired subset. We have solved the program Eq. 15 analytically, while for general cases, convex additional constraints maintains the convexity of the optimization program, for which efficient numerical algorithms can be adopted [59, 60]. With this approach, the minimum change principle may be extended to quantum combs [61, 62], quantum supermaps [63, 64], and quantum Bayesian networks [65, 19, 66, 67, 68], offering new belief update rules for them. The tools introduced in this work may also pave the way towards a fully quantum generalisation of entropy production and fluctuation theorems [69, 70].
Acknowledgments.—G. B. and V. S. acknowledge support from the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; from the National Research Foundation, Singapore, through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009); and from the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). F. B. acknowledges support from MEXT Quantum Leap Flagship Program (MEXT QLEAP) Grant No.  JPMXS0120319794, from MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe,” No. 21H05183, and from JSPS KAKENHI Grant No. 23K03230. G. B. acknowledges support from the Start-up Fund (Grant No. G0101000274) from The Hong Kong University of Science and Technology (Guangzhou).
@@ Appendix A Proof of Theorem 1
We use the definition of $Q_{\rm fwd}$ and $Q_{\rm rev}$ in Eqs. 11 and 12. Since we are first proving the existence of a local optimal solution of Eq. 15 in the form Eq. 16, it suffices to restrict $C_{\mathcal{R}}$ to be full-rank, and then $Q_{\rm rev}$ is full-rank. Recall that the variable we optimize over is $\mathcal{R}$, or equivalently the transpose of its Choi operator $C_{\mathcal{R}}^{T}$.
@@
$(\sqrt{\tau}\otimes\mathds{1}_{A})\Delta(\sqrt{\tau}\otimes\mathds{1}_{A})+\Lambda^{T}\otimes\mathds{1}_{A}$ =0=0= 0 (42)
$\mathrm{Tr}_{A}[C_{\mathcal{R}}]-\mathds{1}_{B}$ =0=0= 0 (43)
@@
where $\Delta_{B}:=-\tau^{-1/2}\Lambda^{T}\tau^{-1/2}$. $\Delta_{B}>0$ since $\Delta>0$. By Eq. 43, one gets $\mathds{1}_{B}=\mathrm{Tr}_{A}[C_{\mathcal{R}}^{T}]$. Substituting Eqs. 39 and 44 into this, one gets
@@
where $D:=\sqrt{\tau}\left(\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau}\right)^{-1/2}$. The above Choi operator corresponds to the map in Eq. 16. To see this, assuming the Kraus decomposition of $\mathcal{E}$ is $\mathcal{E}(\rho)=\sum_{k}K_{k}\rho K_{k}^{\dagger}$, one computes $C_{\mathcal{R}}^{T}$ from Eq. 16 by definition of the Choi operator $C_{\mathcal{R}}=\sum_{ij}|i\rangle\!\!\!\;\langle j|\otimes\mathcal{R}(|i\rangle\!\!\!\;\langle j|)$ of the map $\mathcal{R}(\sigma):=\sqrt{\gamma}\,\mathcal{E}^{\dagger}(D\sigma D^{\dagger})\sqrt{\gamma}$ as
@@
$=\sum_{k}\left((\mathds{1}_{B}\otimes\sqrt{\gamma})\left(\sum_{ij}|i\rangle\!\!\!\;\langle j|\otimes\ K_{k}^{\dagger}D|i\rangle\!\!\!\;\langle j|D^{\dagger}K_{k}\right)(\mathds{1}_{B}\otimes\sqrt{\gamma})\right)^{T}\,.$ (60)
Using the identity $\sum_{ij}|i\rangle\!\!\!\;\langle j|\otimes X|i\rangle\!\!\!\;\langle j|Y=\sum_{ij}X^{T}|i\rangle\!\!\!\;\langle j|Y^{T}\otimes|i\rangle\!\!\!\;\langle j|$ for operators $X$ and $Y$ on ${\mathcal{H}}_{B}$, one has
$C_{\mathcal{R}}^{T}$ $=\sum_{k}\left((\mathds{1}_{B}\otimes\sqrt{\gamma})\left(\sum_{ij}(K_{k}^{\dagger}D)^{T}|i\rangle\!\!\!\;\langle j|(D^{\dagger}K_{k})^{T}\otimes|i\rangle\!\!\!\;\langle j|\right)(\mathds{1}_{B}\otimes\sqrt{\gamma})\right)^{T}$ (61)
@@
By Eq. 26, $\Delta=\sqrt{\rho}\left(\sqrt{\rho}\,\sigma\sqrt{\rho}\right)^{-1/2}\sqrt{\rho}=\sqrt{\rho}M^{-1}\sqrt{\rho}$. Using Eqs. 36 and 69, its differential is
@@
Combining this with Eq. 68, we get the second-order derivative of $f$